OpenStudy (anonymous):

suppose f(x) = 2e^(3x-4). Find a formula for f^-1(x), the inverse function

OpenStudy (campbell_st):

let \[y = 2e^{3x - 4}\] swap the x and y \[x = 2e^{3y -4}\] now may y the subject \[\frac{x}{2} = e^{3y - 4}\] take the base e log of both sides \[\ln \frac{x}{2} = 3y - 4\] add 4 to both sides \[\ln (\frac{x}{2}) + 4 = 3y\] divide by 3 \[y = \frac{1}{3} [\ln(\frac{x}{2}) +4]\]

OpenStudy (campbell_st):

then \[f^{-1} = \frac{1}{3} (\ln(\frac{x}{2}) + 4)\]